OpenStudy (anonymous):

2.Expand (4x – 3y)4 using Pascal’s Triangle

OpenStudy (anonymous):

(4x – 3y)4= 4*(4x – 3y)= 4*4x - 4*3y = 16x - 12y

OpenStudy (anonymous):

Kumar has read the 4 on the end as "times 4" but I think you must mean "to the power of 4" pascal's triangle is 1 1 1 1 2 1 1 3 3 1 1 4 6 4 1 those numbers are the coefficients, & what's inside the brackets gets raised to powers, and add all the terms together: \[(4x)^4.(-3y)^0 + 4(4x)^3.(-3y)^1 + 6(4x)^2.(-3y)^2. + 4(4x)^1.(-3y)^3 + (4x)^0.(-3y)^4\] work out all the powers of the numbers INSIDE the brackets, multiply those by the coefficients from pascal's triangle & be careful with your negatives. Anything to the power of 0 is equal to 1. Notice that the powers of the two brackets always add to 4.